# Entangling Quantum Memories at Channel Capacity

### v1

@@ Abstract
We plot the performance of the link for different values of the half-channel loss $(\sqrt{\eta})$ for various $N$-qubit memory registers (varying colors) utilizing GKP qudits of appropriate dimension $(d=2^{N})$ in Fig. 4 for both the sq-GKP (a) and hex-GKP (b) encodings. The performance of infinite squeezed states (solid; $\sigma_{i}^{2}\rightarrow 0$) is shown as a benchmark; finite squeezed states for 10 dB (dot-dashed; $+$ markers) and 5 dB (dashed; $\circ$ markers) of squeezing are shown for comparison. We compare these curves with the repeater-less channel capacity of $-\log_{2}(1-\sqrt{\eta})$ (black; dotted). For $\eta\rightarrow 0$ dB, we note that $\mathcal{R}(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})\rightarrow N$. Thus by increasing the size of the memory registers, the rate scaling of $-\log_{2}(1-\sqrt{\eta})$ for $\eta\rightarrow 1$ can be approached (with an offset of $\log_{2}e$ similar to Ref. [6]) by the proposed link. Realistic implementations of the high-rate links will be affected by the CSUM gate implementation as well as the finite squeezing of the GKP states - we have analyzed some examples in the Methods.
@@
Figure 4: Performance of the qudit-swap assisted link for $N$-qubit memory registers (varying colors) utilizing (a) square lattice and (b) hexagonal latticle GKP qudit encodings $(d=2^{N})$. For both lattice choices, we analyze the performance of protocol for infinite squeezed states (solid), 5 dB (dashed; $\circ$ markers) and 10 dB (dot-dashed; $+$ markers) squeezed states with the channel capacity of $C(\eta)$.
@@
We also find that rate of entanglement with large dimensional encodings ($d\rightarrow\infty$) with infinitely squeezed GKP qudit states ($\sigma_{i}^{2}\rightarrow 0$) approach the repeater-less bound $Q_{2}(\sqrt{\eta})=-\log_{2}(1-\sqrt{\eta})$ ebits per channel use for the entanglement swap in the low loss limit ($\eta\rightarrow 1$). Our detailed proof is included in the Methods - we outline the highlights of the proof here — In the limit of small state variance it is sufficient to lower bound the shift error probabilities $P(X^{k},\sigma^{2})=P(Z^{k},\sigma^{2})$ by dropping the summation over the lattice positions (i.e. $j$ index) in Eq. (18) - we may label this lower bound $\tilde{P}_{\mathrm{shift}}({k};\sigma^{2})$. In the regime of interest, it is also sufficient to drop larger shifts by considering $\tilde{P}_{\mathrm{shift}}({k};\sigma^{2})\approx 0$ for $|k|>1$, i.e. the probability of incurring larger (greater than one shift) logical errors is negligible. Carrying out a Taylor expansion of the error function around $d\rightarrow\infty$ (since we expect the system size requirements to increase) and $\eta\rightarrow 1$, gives us the simplified expressions
@@
$\begin{split}&\tilde{P}^{(\mathrm{hex})}_{\mathrm{shift}}({0};\varepsilon)=1-\frac{\sqrt{2d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\pi}\\ &\tilde{P}^{(\mathrm{hex})}_{\mathrm{shift}}(\pm 1;\varepsilon)=\frac{\sqrt{d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\sqrt{2}\pi}\end{split}$ (21)
@@
where $\varepsilon=1-\sqrt{\eta}$. We may make the ansatz that $2d\varepsilon=\xi$ where $\xi\in\mathbb{R}$ is an appropriately chosen constant that captures the required qudit dimensionality requirements.
@@
The hashing bound with all the approximations stated above (which we may label $I_{\mathrm{LB}}(\xi)$) is lower than the hashing bound $I(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})$. The matter of choosing an optimal $\xi$ to obtain the best rate scaling is equivalent to minimizing its difference from the channel capacity, i.e. we seek to minimize $Q_{2}(\sqrt{\eta})-I_{\mathrm{LB}}(\xi)$. This may be performed numerically; choosing $\xi_{\mathrm{opt}}\approx 1.642$ gives us a separation of $\sim 1.06$ ebits per channel use between $I_{\mathrm{LB}}(\xi)$ and the repeaterless bound $Q_{2}(\sqrt{\eta})$.
@@
In this case, an ansatz of $\sqrt{3}d\varepsilon=\xi^{\prime}$, allows us to come up with a similar expression for $I_{\mathrm{LB}}^{(\mathrm{hex})}(\xi^{\prime})$ and to optimize the rate lower bound separation from $Q_{2}(\sqrt{\eta})$. By choosing $\xi^{\prime}_{\mathrm{opt}}\approx 1.422$ we obtain a separation of $\sim 0.85$ ebits per channel – this matches the expectation that $\xi_{\mathrm{opt}}/\xi^{\prime}_{\mathrm{opt}}=\sqrt{3}/2$, i.e. the qudit dimesionality vs. loss scaling is proportional to the improvement in the error probability between the lattice choices.
@@
Figure 5: Low-loss limit performance of the qudit-swap assisted link (a) Comparison of varying $N$ (varying colors) utilizing sq-GKP (solid lines) and hex-GKP (dot-dashed) qudit encodings $(d=2^{N})$. We compare the finite squeezing, finite encoding size performance with calculated asymptotic rate scaling (red: sq-GKP; blue: hex-GKP) and the repeater-less channel capacity of $Q_{2}(\eta)=-\log_{2}(1-\sqrt{\eta})$ (black, dotted). (b) Comparison of the asymptotic expressions and the repeater-less bound showing a finite separation in the extremely low-loss regime i.e. $\varepsilon\rightarrow 0;\eta\rightarrow 1$.
Possible Extensions
The CSUM gate protocol proposed in this work is a key element for hybrid quantum logic between memory registers and GKP qudits – the applications proposed in Ref. [1] can naturally be extended to higher dimensional encoding. It is pertinent to note that protocols with higher-dimensional encodings are expected to be more sensitive to any imperfections in the cavity-photon interface i.e. lack of strong coupling, atomic and cavity loss, as well as the finite squeezing level of the GKP states. Further investigations of the underlying trade-offs and thresholds in the protocol performance are left for future works.
@@
We have presented a method for generating entanglement between a register of $N$-quantum memories and $d=2^{N}$ dimensional GKP qudit states, assisted by the interaction of bosonic modes with cavity QED systems. We demonstrated how the resource states generated by our protocol could be utilized for the implementation of high rate quantum links especially in the low-loss regime of operation. Furthermore, we show that the hashing bound of the final state (which is a lower bound to its distillable entanglement) scales as the repeaterless bound $C(\eta)$ for entanglement swapping in the low-loss regime with commensurately chosen $N$.State-of-the-art systems have demonstrated internal cooperativities of the order $\sim 100$. Improved cavity designs with larger cooperativities are being actively pursued in the various physical platforms that could support the necessary interactions. Coupled with efforts for large scale integration of physical quantum memory systems [20], the protocols suggested here could be achievable in the near term. We have also discussed a few potential applications of the generalized interaction; these are pedagogically interesting for various tasks in higher-dimensional quantum resource state generation for fault tolerant measurement based quantum computing. Our findings demonstrate a pathway to achieving channel capacity saturating quantum communication links.
@@ Acknowledgements
S.G. has outside interests in SensorQ Technologies Incorporated and Guha, LLC. These interests have been disclosed to UA and reviewed in accordance with their conflict of interest policies, with any conflicts of interest to be managed accordingly.
@@ Methods
Similar to how the standard Pauli operators for qubits, we can define generalized higher dimensional logical operators. The operators can be expressed either in terms of Gell-Mann matrices or using the Weyl operators [21] - we shall use the latter to denote the system operations. The standard Weyl operators for a $d$-dimensional qudit Hilbert space is defined as the set $\mathcal{W}=\{W^{(n,m)}|n,m=0,\ldots,d-1\}$, where,
@@
Maximally entangled states of two $d$-dimensional qudit systems (labelled by their subscripts $1$ and $2$) are denoted by |$\ket{\Psi_{k,l}}_{1,2}$, which are defined as
@@
$\hat{a}_{\text{in}}(\omega)$ $=r(\omega)\hat{a}_{\text{out}}(\omega)+l_{C}(\omega)\hat{a}_{l}(\omega)+l_{A}(\omega)\hat{a}_{\gamma}(\omega)$ (28)
@@
$r(\omega)$ $=1-\frac{2\zeta i}{\left(1-2i{\Delta_{c}}/{\kappa}+{C}/{(1-2i\Delta_{a}/\gamma_{m})}\right)}$ (29a)
$l_{C}(\omega)$ $=-\frac{2\sqrt{\zeta\,(1-\zeta)}}{\left(1-2i{\Delta_{c}}/{\kappa}+{C}/{(1-2i\Delta_{a}/\gamma_{m})}\right)}$ (29b)
@@
In any interaction, modes represented by $\hat{a}_{\gamma}$ and $\hat{a}_{l}$ will be lost since they represent scattered field operators - hence it is appropriate to call $r(\omega)$ the reflectivity, $l_{C}(\omega)$ the cavity loss coefficient and $l_{A}(\omega)$ the atomic loss coefficient. Readers may note that the virtual excitation condition (i.e. $\langle\sigma_{z}\rangle=-1$) is not true [17] when $\Delta_{a}\rightarrow 0$; however the cavity-atom system still imparts a memory dependent phase on the travelling modes [15].
@@
|$\ket{\alpha}_{B_{j}}$ $\equiv\bigotimes_{\omega}\ket{\alpha_{j}f_{j}(\omega)d\omega}_{B_{j}}$ (30)
@@
|$\ket{\phi}_{M_{j},B_{j}}=\sum_{s=\{0,1\}}c_{j,s}\ket{s}_{M,j}\bigotimes_{\omega}\ket{r^{(s)}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{B_{j}}\ket{l_{C}^{(s)}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{C,j}}\ket{l_{A}^{(s)}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{A,j}}$ (31)
@@
Thus starting with initial memory state of |$\ket{\Psi}_{\bm{M}}=\sum_{\bm{m}}c_{\bm{m}}\ket{\bm{m}}_{\bm{M}}$ yields the final state |$\ket{\Phi}_{\bm{M},\bm{B}}=\sum_{\bm{m}}c_{\bm{m}}\ket{\phi_{\bm{m}}}_{\bm{M},\bm{B}}$, where upon tracing out the cavity and atomic loss modes we obtain the state
[…]
@@
$\beta_{\bm{m}}$ $=\bigotimes_{j=1}^{N}\bigotimes_{\omega}\ket{r^{(m_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{B_{j}}$ (34a)
$g_{\bm{m},\bm{m}^{\prime}}$ $=c_{\bm{m}}c^{*}_{\bm{m}^{\prime}}\left(\delta_{\bm{m}^{\prime},\bm{m}^{\prime}}+(1-\delta_{\bm{m}^{\prime},\bm{m}^{\prime}})\lambda_{\bm{m}^{\prime},\bm{m}^{\prime}}\right)$ (34b)
$\lambda_{\bm{m}^{\prime},\bm{m}^{\prime}}$ $=\bigotimes_{j=1}^{N}\bigotimes_{\omega}\braket{l_{C}^{(m_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}{l_{C}^{(m^{\prime}_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{C,j}}\times\braket{l_{A}^{(m_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}{l_{A}^{(m^{\prime}_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{A,j}}$ (34c)
@@
$g_{\bm{m},\bm{m}^{\prime}}$ $=c_{\bm{m}}c^{*}_{\bm{m}^{\prime}}\left(\delta_{\bm{m}^{\prime},\bm{m}^{\prime}}+(1-\delta_{\bm{m}^{\prime},\bm{m}^{\prime}})|\lambda_{\bm{m}^{\prime},\bm{m}^{\prime}}|^{2}\right)$ (35)
@@
$=$ $\frac{\sqrt{2\pi/d}}{2}(2\bm{m}-(d-1))=(\bm{m}-(d-1)/2)\times\sqrt{2\pi/d}$ (36d)
Here, we use the binary representation of $\bm{m}=[m_{1}m_{2}\ldots m_{N}]_{2}=\sum_{k=1}^{N}m_{k}\times 2^{N-k}$. Additionally in Eq. (36c), we first use the equivalence between the functions $g(m_{k})=(m_{k}-\bar{m}_{k})$ and $f(m_{k})=(-1)^{m_{k}+1}$ for $m_{k}=\{0,1\}$, where $\bar{m}_{k}$ is the complement of ${m}_{k}$. We use $\bar{\bm{m}}$ to represent the decimal representation of an integer obtained by bitwise complement of the bits of $\bm{m}$, i.e., $\bar{\bm{m}}=[\bar{m}_{1}\bar{m}_{2}\ldots\bar{m}_{N}]_{2}$. In Eq. (36d), we use the identity $\bar{\bm{m}}=(2^{N}-1)-\bm{m}=(d-1)-\bm{m}$ to obtain the final expression. We list the necessary coherent amplitudes, and pre-displacements for the various gates that can be implemented by the proposed interactions in Table 1
@@
Table 1: Required coherent amplitudes for relected pulses and pre-displacement to implement various memory register to GKP ($\bm{M}\rightarrow G$) qudit gates.
.3 Quality of Entanglement Swap based on Dual Homodyne Measurement
To obtain the exact effect of pre-amplification + pure loss we have to calculate the action of the Kraus operators on the complete quantum state. However, for GKP qubits this process is easier said than done due to the complicated state description. A few alternatives (shortcuts) to this process can be implemented -
@@
    Treat the effect of pre-amplification + pure loss as an effective modifier to the individual peak variance for the GKP state while preserving initial state purity.
@@
    ‘Twirl’ the combined pre-amplification + pure loss channel to express its effect as effective Pauli errors on the state.
We expect that the former approach will overestimate the quality of the final state since it treats the effect of loss coherently - i.e. the final state is pure. The latter will give us an achievable lower bound since twirling the state ignores any residual coherences in the state - however it allows for a simplified calculation of the hashing rate as we shall show subsequently.
Assuming we start with the ideal memory-GKP maximally entangled state,

### v2

@@ Abstract
We plot the performance of the link for different values of the half-channel loss $(\sqrt{\eta})$ for various $N$-qubit memory registers (varying colors) utilizing GKP qudits of appropriate dimension $(d=2^{N})$ in Fig. 4 (for pre-amplification of the transmitted state) and Fig. 5 (for CC-amplification of the measurement results) for both the sq-GKP (a) and hex-GKP (b) encodings. The performance of infinite squeezed states (solid; $\sigma_{i}^{2}\rightarrow 0$) is shown as a benchmark; finite squeezed states for 10 dB (dot-dashed; $+$ markers) and 5 dB (dashed; $\circ$ markers) of squeezing are shown for comparison. We compare these curves with the repeater-less channel capacity of $-\log_{2}(1-\sqrt{\eta})$ (black; dotted). In Fig. 4 (5) we have included the performance of the CC-amplification (pre-amplification) protocol as semi-transparent color matched dot-dashed curves for various values $N$.
For $\eta\rightarrow 0$ dB, we note that $\mathcal{R}(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})\rightarrow N$. Thus by increasing the size of the memory registers, the rate scaling of $-\log_{2}(1-\sqrt{\eta})$ for $\eta\rightarrow 1$ can be approached (with an offset of $\log_{2}e$ similar to Ref. [7]) by the proposed link. Furthermore the CC-amplification protocol (which adds lower variance noise), outperforms the pre-amplification for all values of half channel loss upto 3dB (as noted by the transformation rules earlier). Realistic implementations of the high-rate links will be affected by the CSUM gate implementation as well as the finite squeezing of the GKP states - we have analyzed some examples in the Methods.
@@
Figure 4: Performance of the qudit-swap assisted link for $N$-qubit memory registers (varying colors) utilizing (a) square lattice and (b) hexagonal lattice GKP qudit encodings $(d=2^{N})$ assuming pre-amplification of the transmitted qudits. For both lattice choices, we analyze the performance of protocol for infinite squeezed states (solid), 5 dB (dashed; $\circ$ markers) and 10 dB (dot-dashed; $+$ markers) squeezed states with the channel capacity of $C(\eta)$.
Figure 5: Performance of the qudit-swap assisted link for $N$-qubit memory registers (varying colors) utilizing (a) square lattice and (b) hexagonal lattice GKP qudit encodings $(d=2^{N})$ assuming the CC-amplification of dual-homodyne measurement outcomes. For both lattice choices, we analyze the performance of protocol for infinite squeezed states (solid), 5 dB (dashed; $\circ$ markers) and 10 dB (dot-dashed; $+$ markers) squeezed states with the channel capacity of $C(\eta)$.
@@
We also find that rate of entanglement with large dimensional encodings ($d\rightarrow\infty$) with infinitely squeezed GKP qudit states ($\sigma_{i}^{2}\rightarrow 0$) approach the repeater-less bound $Q_{2}(\sqrt{\eta})=-\log_{2}(1-\sqrt{\eta})$ ebits per channel use for the entanglement swap in the low loss limit ($\eta\rightarrow 1$). Our detailed proof is included in the Methods - we outline the highlights of the proof here — In the limit of small state variance it is sufficient to lower bound the shift error probabilities $P(X^{k},\sigma^{2})=P(Z^{k},\sigma^{2})$ by dropping the summation over the lattice positions (i.e. $j$ index) in Eq. (19) - we may label this lower bound $\tilde{P}_{\mathrm{shift}}({k};\sigma^{2})$. In the regime of interest, it is also sufficient to drop larger shifts by considering $\tilde{P}_{\mathrm{shift}}({k};\sigma^{2})\approx 0$ for $|k|>1$, i.e. the probability of incurring larger (greater than one shift) logical errors is negligible. Carrying out a Taylor expansion of the probability of shift error around $d\rightarrow\infty$ (since we expect the system size requirements to increase) and $\eta\rightarrow 1$, gives us the simplified expressions
@@
$\begin{split}&\tilde{P}^{(\mathrm{sq})}_{\mathrm{shift}}({0};\varepsilon)=1-\frac{\sqrt{2d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\pi}\\ &\tilde{P}^{(\mathrm{sq})}_{\mathrm{shift}}(\pm 1;\varepsilon)=\frac{\sqrt{d\varepsilon}\exp\left(-\pi/(2d\varepsilon)\right)}{\sqrt{2}\pi}\end{split}$ (22)
@@
where $\varepsilon=1-\sqrt{\eta}$ for the pre-amplification protocol. We may make the ansatz that $2d\varepsilon=\xi$ where $\xi\in\mathbb{R}$ is an appropriately chosen constant that captures the required qudit dimensionality requirements.
@@
The hashing bound with all the approximations stated above (which we may label $I_{\mathrm{LB}}(\xi)$) is lower than the hashing bound $I(\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}})$. The matter of choosing an optimal $\xi$ to obtain the best rate scaling is equivalent to minimizing its difference from the channel capacity, i.e. we seek to minimize $Q_{2}(\sqrt{\eta})-I_{\mathrm{LB}}(\xi)$. This may be performed numerically; choosing $\xi_{\mathrm{opt}}\approx 1.642$ gives us a separation of $\sim 1.06$ ebits per channel use between $I_{\mathrm{LB}}(\xi)$ and the repeaterless bound $Q_{2}(\sqrt{\eta})$. A similar ansatz can be made for the CC-amplification protocol where $\sqrt{\eta}=1-\varepsilon\Rightarrow(1-\sqrt{\eta})/(2\sqrt{\eta})\approx\varepsilon/2$. Then by choosing $d\varepsilon=\xi_{\mathrm{opt}}$ we obtain the same scaling as that of the pre-amplification protocol.
@@
In this case, an ansatz of $\sqrt{3}d\varepsilon=\xi^{\prime}$, allows us to come up with a similar expression for $I_{\mathrm{LB}}^{(\mathrm{hex})}(\xi^{\prime})$ and to optimize the rate lower bound separation from $Q_{2}(\sqrt{\eta})$. By choosing $\xi^{\prime}_{\mathrm{opt}}\approx 1.422$ we obtain a separation of $\sim 0.85$ ebits per channel – this matches the expectation that $\xi_{\mathrm{opt}}/\xi^{\prime}_{\mathrm{opt}}=\sqrt{3}/2$, i.e. the qudit dimesionality vs. loss scaling is proportional to the improvement in the error probability between the lattice choices. Similarly for the CC-amplification, we may make the ansatz $\sqrt{3}d\varepsilon/2=\xi^{\prime}_{\mathrm{opt}}$.
@@
Figure 6: Low-loss limit performance of the qudit-swap assisted link (a) Comparison of varying $N$ (varying colors) utilizing sq-GKP (solid lines) and hex-GKP (dot-dashed) qudit encodings $(d=2^{N})$ for the pre-amplification protocol. We compare the finite squeezing, finite encoding size performance with calculated asymptotic rate scaling (red: sq-GKP; blue: hex-GKP) and the repeater-less channel capacity of $Q_{2}(\eta)=-\log_{2}(1-\sqrt{\eta})$ (black, dotted). (b) Comparison of the asymptotic expressions (for either protocol) and the repeater-less bound showing a finite separation in the extremely low-loss regime i.e. $\varepsilon\rightarrow 0;\eta\rightarrow 1$.
@@
We have presented a method for generating entanglement between a register of $N$-quantum memories and $d=2^{N}$ dimensional GKP qudit states, assisted by the interaction of bosonic modes with cavity QED systems. We demonstrated how the resource states generated by our protocol could be utilized for the implementation of high rate quantum links especially in the low-loss regime of operation, thereby highlighting a pathway to achieving channel capacity saturating quantum communication links.
Furthermore, we show that the achievable entanglement generation rate scales as the repeaterless bound $C(\eta)$ for entanglement swapping in the low-loss regime with commensurately chosen $N$. State-of-the-art systems have demonstrated internal cooperativities of the order $\sim 100$. Improved cavity designs with larger cooperativities are being actively pursued in the various physical platforms that could support the necessary interactions. Coupled with efforts for large scale integration of physical quantum memory systems [23], the protocols suggested here could be achievable in the near term.
The CSUM gate protocol proposed in this work is a key element for hybrid quantum logic between memory registers and GKP qudits – the applications proposed in Ref. [2] can naturally be extended to high dimensional encodings. It is pertinent to note that protocols with higher-dimensional encodings are expected to be more sensitive to any imperfections in the cavity-photon interface. Further investigations of the underlying trade-offs and thresholds in the protocol performance are left for future works.
@@ Acknowledgements
Conflicts of Interest
S.G. has outside interests in SensorQ Technologies Incorporated and Guha, LLC, and holds shares in Quantum Network Technologies (Qunett). These interests have been disclosed to UMD and reviewed in accordance with its conflict of interest (CoI) policies, with any resulting CoI to be managed accordingly.
@@ Methods
Similar to how the standard Pauli operators are defined for qubits, we can define generalized higher dimensional logical operators. The operators can be expressed either in terms of Gell-Mann matrices or using the Weyl operators [24] - we shall use the latter to denote the system operations. The standard Weyl operator set for a $d$-dimensional qudit Hilbert space is defined as the set $\mathcal{W}=\{W^{(n,m)}|n,m=0,\ldots,d-1\}$, where,
@@
Maximally entangled states of two $d$-dimensional qudit systems (labelled by their subscripts $1$ and $2$) are denoted by |$\ket{\Psi_{k,l}}_{1,2}$, which are defined as,
@@
$\hat{a}_{\text{in}}(\omega)$ $=r(\omega)\hat{a}_{\text{out}}(\omega)+l_{C}(\omega)\hat{a}_{l}(\omega)+l_{A}(\omega)\hat{a}_{\gamma}(\omega),$ (29)
@@
$r(\omega)$ $=1-\frac{2\zeta i}{\left(1-2i{\Delta_{c}}/{\kappa}+{C}/{(1-2i\Delta_{a}/\gamma_{m})}\right)};$ (30a)
$l_{C}(\omega)$ $=-\frac{2\sqrt{\zeta\,(1-\zeta)}}{\left(1-2i{\Delta_{c}}/{\kappa}+{C}/{(1-2i\Delta_{a}/\gamma_{m})}\right)};$ (30b)
@@
In any interaction, modes represented by $\hat{a}_{\gamma}$ and $\hat{a}_{l}$ will be lost since they represent scattered field operators - hence it is appropriate to call $r(\omega)$ the reflectivity, $l_{C}(\omega)$ the cavity loss coefficient and $l_{A}(\omega)$ the atomic loss coefficient. An improved approach with similar conclusions is shown in a recent work [18].
@@
|$\ket{\alpha}_{B_{j}}$ $\equiv\bigotimes_{\omega}\ket{\alpha_{j}f_{j}(\omega)d\omega}_{B_{j}}.$ (31)
@@
|$\ket{\phi}_{M_{j},B_{j}}=\sum_{s=\{0,1\}}c_{j,s}\ket{s}_{M,j}\bigotimes_{\omega}\ket{r^{(s)}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{B_{j}}\ket{l_{C}^{(s)}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{C,j}}\ket{l_{A}^{(s)}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{A,j}},$ (32)
@@
Thus starting with initial memory state of |$\ket{\Psi}_{\bm{M}}=\sum_{\bm{m}}c_{\bm{m}}\ket{\bm{m}}_{\bm{M}}$ yields the final state |$\ket{\Phi}_{\bm{M},\bm{B}}=\sum_{\bm{m}}c_{\bm{m}}\ket{\phi_{\bm{m}}}_{\bm{M},\bm{B}}$, where upon tracing out the cavity and atomic loss modes we obtain the state,
$\rho_{\bm{M},\bm{B}}=\frac{1}{d}\sum^{2^{N}}_{\bm{m},\bm{m}^{\prime}=1}g_{\bm{m},\bm{m}^{\prime}}\;\ket{\bm{m}}\!\!\bra{\bm{m}^{\prime}}_{\bm{M}}\otimes\ket{\beta_{\bm{m}}}\!\!\bra{\beta_{\bm{m}^{\prime}}}_{\bm{B}},$ (34)
@@
$\beta_{\bm{m}}$ $=\bigotimes_{j=1}^{N}\bigotimes_{\omega}\ket{r^{(m_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{B_{j}};$ (35a)
$g_{\bm{m},\bm{m}^{\prime}}$ $=c_{\bm{m}}c^{*}_{\bm{m}^{\prime}}\left(\delta_{\bm{m}^{\prime},\bm{m}^{\prime}}+(1-\delta_{\bm{m}^{\prime},\bm{m}^{\prime}})\lambda_{\bm{m}^{\prime},\bm{m}^{\prime}}\right);$ (35b)
$\lambda_{\bm{m}^{\prime},\bm{m}^{\prime}}$ $=\bigotimes_{j=1}^{N}\bigotimes_{\omega}\braket{l_{C}^{(m_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}{l_{C}^{(m^{\prime}_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{C,j}}\times\braket{l_{A}^{(m_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}{l_{A}^{(m^{\prime}_{j})}(\omega)\alpha_{j}f_{j}(\omega)d\omega}_{L_{A,j}}.$ (35c)
@@
$g_{\bm{m},\bm{m}^{\prime}}$ $=c_{\bm{m}}c^{*}_{\bm{m}^{\prime}}\left(\delta_{\bm{m}^{\prime},\bm{m}^{\prime}}+(1-\delta_{\bm{m}^{\prime},\bm{m}^{\prime}})|\lambda_{\bm{m}^{\prime},\bm{m}^{\prime}}|^{2}\right).$ (36)
@@
$=$ $\frac{\sqrt{2\pi/d}}{2}(2\bm{m}-(d-1))=(\bm{m}-(d-1)/2)\times\sqrt{2\pi/d}.$ (37d)
Here, we use the binary representation of $\bm{m}=[m_{1}m_{2}\ldots m_{N}]_{2}=\sum_{k=1}^{N}m_{k}\times 2^{N-k}$. Additionally in Eq. (37c), we first use the equivalence between the functions $g(m_{k})=(m_{k}-\bar{m}_{k})$ and $f(m_{k})=(-1)^{m_{k}+1}$ for $m_{k}=\{0,1\}$, where $\bar{m}_{k}$ is the complement of ${m}_{k}$. We use $\bar{\bm{m}}$ to represent the decimal representation of an integer obtained by bitwise complement of the bits of $\bm{m}$, i.e., $\bar{\bm{m}}=[\bar{m}_{1}\bar{m}_{2}\ldots\bar{m}_{N}]_{2}$. In Eq. (37d), we use the identity $\bar{\bm{m}}=(2^{N}-1)-\bm{m}=(d-1)-\bm{m}$ to obtain the final expression. We list the necessary coherent amplitudes, and pre-displacements for the various gates that can be implemented by the proposed interactions in Table 1.
@@
Table 1: Required coherent amplitudes for reflected pulses and pre-displacement to implement various memory register to GKP ($\bm{M}\rightarrow G$) qudit gates.
.3 Bounding the Quality of Distributed Entanglement
A faithful analysis of the proposed quantum link would require a complete evolution of the transmitted GKP qudit states through the pure loss channel (with possible pre-amplification). However, given the non-Gaussian state description of the underlying states, this process is complicated. Analyses by prior studies [25, 5] consider a ‘twirled’ error approach - we present an argument why this approach presents a conservative estimate to the final state quality.
We implicitly assume that the pure loss channel is augmented by amplification (either pre-amplification of state of post-amplification of the measurement outcomes), thereby effectively converting it into a random displacement channel, $\mathcal{N}_{B_{2}}[\sigma^{2}]$. Given an initial pure approximate GKP qudit state |$\ket{\chi}_{G}$ , the channel output is described by
$\mathcal{N}_{B_{2}}[\sigma^{2}](\rho_{G})=\frac{1}{\pi\sigma^{2}}\int d^{2}\alpha\,e^{-|\alpha|^{2}/\sigma^{2}}D(\alpha)\ket{\chi}\!\!\bra{\chi}_{G}D^{\dagger}(\alpha).$ (38)
Rather than evaluating the continuous integral over phase space, a few simplified approaches are possible -
@@
    We may treat the effect of random displacement as an effective modifier to the individual peak variance for the GKP state while preserving initial state purity. Given the initial state |$\ket{\chi}_{G}$ expressed in the approximate GKP basis, Eq. (4) with some initial variance $\sigma^{2}_{i}$, the final state is a pure state |$\ket{\chi^{\prime}}_{G}$ with a modified variance $\sigma_{i}^{2}+g(\sigma^{2})$, where the modifier $g(\sigma^{2})$ is a function of the random noise parameter.
@@
    We may ‘twirl’ the channel and approximate the continuous random displacements as discretized Pauli errors with a weighted error probability. This is described by
    $\frac{1}{\pi\sigma^{2}}\int d^{2}\alpha\,e^{-|\alpha|^{2}/\sigma^{2}}D(\alpha)\ket{\chi}\!\!\bra{\chi}_{G}D^{\dagger}(\alpha)\rightarrow$ $\sum_{k_{1},k_{2}}\int_{\mathbb{L}(k_{1},k_{2})}d^{2}\alpha\,e^{-|\alpha|^{2}/\sigma^{2}}D(\alpha)\ket{\chi}\!\!\bra{\chi}_{G}D^{\dagger}(\alpha)$
    $\approx\sum_{k_{1},k_{2}}P(X^{k_{1}})P(Z^{k_{2}})\hat{X}^{k_{1}}\hat{Z}^{k_{2}}\ket{\chi}\!\!\bra{\chi}_{G}\hat{Z}^{k_{2}}\hat{X}^{k_{1}},$ (39)
    where $\mathbb{L}(k_{1},k_{2})$ is the lattice tile corresponding to $k_{1}$ shifts along the logical Pauli $\hat{X}$ and $k_{2}$ shifts along the logical Pauli $\hat{Z}$ directions for the qudit on phase space. The probabilities $P(X^{k_{1}})$ and $P(Z^{k_{2}})$ correspond to the integrated probability over this lattice tile, while $\hat{X}^{k_{1}}=W_{d}^{(0,k_{1})}$ and $\hat{Z}^{k_{2}}=W_{d}^{(k_{2},0)}$ are the qudit basis traversal operators.
Assuming both parties start with the ideal memory-GKP maximally entangled state,